# Differentially Private Two-Stage Gradient Descent for Instrumental Variable Regression

### v1

@@ Abstract
We study instrumental variable regression (IVaR) under differential privacy constraints. Classical IVaR methods (like two-stage least squares regression) rely on solving moment equations that directly use sensitive covariates and instruments, creating significant risks of privacy leakage and posing challenges in designing algorithms that are both statistically efficient and differentially private. We propose a noisy two-state gradient descent algorithm that ensures $\rho$-zero-concentrated differential privacy by injecting carefully calibrated noise into the gradient updates. Our analysis establishes finite-sample convergence rates for the proposed method, showing that the algorithm achieves consistency while preserving privacy. In particular, we derive precise bounds quantifying the trade-off among privacy parameters, sample size, and iteration-complexity. To the best of our knowledge, this is the first work to provide both privacy guarantees and provable convergence rates for instrumental variable regression in linear models. We further validate our theoretical findings with experiments on both synthetic and real datasets, demonstrating that our method offers practical accuracy-privacy trade-offs.
@@ 1 Introduction
Instrumental variable regression (IVaR) is a key tool in causal inference, designed to recover structural parameters when standard estimators fail due to endogeneity. In many observational settings, covariates are influenced by unobserved confounders, causing naive methods (such as the ordinary least squares (OLS) in the context of linear regression) to produce biased and inconsistent estimates. IVaR circumvents this by leveraging instruments, which are variables that are predictive of the endogenous regressors but independent of hidden confounders, to enable consistent estimation of causal effects (Hausman, 2001; Wooldridge, 2010; Angrist and Krueger, 2001). This perspective is increasingly important in machine learning, for example in recommendation systems where user exposure is confounded by prior preferences (Si et al., 2022), or in reinforcement learning where actions and rewards are jointly influenced by unobserved context (Xu et al., 2023). In such settings, IVaR provides a principled way to disentangle causal effects from spurious correlations, enabling more reliable decision making. However, many applications of IVaR involve sensitive data, such as individual health records, financial transactions, or user interactions, where protecting privacy is of paramount importance. In such settings, releasing model estimates or even intermediate statistics can leak information about individuals in the dataset. Differential privacy (DP) (Dwork et al., 2006) provides a mathematically rigorous framework to ensure that an algorithm’s output does not reveal sensitive information about any single data point. Despite the importance of IVaR in causal inference, to the best of our knowledge, there are no prior works addressing the problem of performing IVaR under differential privacy. This gap motivates the central question of this paper:
@@
Our work focuses on answering this question in the context of linear regression models. To situate our contributions, we briefly review existing work on DP methods for OLS regression, with additional discussion in Section 1.1. Several predominant approaches have emerged in the literature: (i) perturbion methods, where the empirical covariance and cross-covariance matrices are privatized before solving the normal equations; (ii) consensus-based methods, including propose-test-release and exponential mechanism approaches, which directly privatize the estimator through carefully designed randomized output rules; and (iii) gradient perturbation methods, where iterative optimization algorithms are made private by clipping gradients and injecting calibrated Gaussian noise. While all three approaches ensure differential privacy, gradient perturbation combined with clipping has been shown to yield the sharpest statistical rates in OLS regression, particularly in high-dimensional and finite-sample regimes (Bassily et al., 2014; Brown et al., 2024).
Given the centrality of IVaR in causal inference, it is natural to explore whether the aforementioned techniques can be adapted to this setting. Unlike OLS, however, IVaR is based on moment conditions involving both covariates and instruments, making it less straightforward to design private algorithms. In particular, sufficient-statistics perturbation and consensus-based methods have not been explored, and their adaptation is non-trivial due to the inherent ill-posedness of IVaR under weak instruments and the sensitivity of the moment equations. Motivated by the success of gradient-based DP methods in OLS, we focus on extending the noisy gradient descent framework to IVaR, carefully analyzing the interplay between contraction rate, privacy guarantees, and sample size. Specifically, we make the following contributions in this work:
  • We introduce DP-2S-GD (Algorithm 1), the first differentially private algorithm for instrumental variable regression, based on noisy gradient descent with gradient clipping.
  • We establish finite-sample convergence rates for DP-2S-GD (Theorem 3.1), explicitly characterizing the trade-off between privacy, contraction rate, and sample size. The main technical challenge is to carefully control the interaction between privacy-induced noise and the contraction of the gradient dynamics across iterations, with the privacy guarantee ensured by Proposition 3.1.
@@ 1.1 Related work
Differential Privacy for Regression. One can group private regression methods into the following broad families. (1) Output/objective perturbation (private empirical risk minimization (ERM)): add noise to the final estimator (output perturbation) or inject a random linear/quadratic term into a strongly convex loss before optimizing (objective perturbation); these one-shot mechanisms give $(\varepsilon,\delta)$-DP guarantees and excess-risk bounds for convex ERM (Chaudhuri et al. (2011); Kifer et al. (2012); Bassily et al. (2014)). Recent refinements, e.g. Redberg et al. (2023), leverage subsampling and tighter accounting to improve accuracy. (2) Sufficient-statistics (matrix) perturbation: release noisy surrogates of ($\mathbf{X}^{\top}\mathbf{X},\mathbf{X}^{\top}\mathbf{y}$) (or related second-moment structures) and then solve the (regularized) normal equations; this route enables OLS-specific inference but can suffer under ill-conditioning because noise is injected at the Gram-matrix level (Dwork et al. (2014); Sheffet (2017)). Further developments in this direction include Bernstein and Sheldon (2019) and  Ferrando and Sheldon (2024). (3) Exponential mechanism: privately selects an output by randomly choosing among candidates with probabilities that grow exponentially with their quality score, with parameters controlling how strongly it favors the higher-scoring options. This mechanism is frequently applied in constructing algorithm to privately select a regression model from a pool of non-private OLS fits on subsets of the data (Ramsay and Chenouri (2021), Cumings-Menon (2022), Amin et al. (2022)). (4) Gradient perturbation (DP-(S)GD): clip per-example (mini-batch or full) gradients and add Gaussian noise at each step, tracking privacy with bounded log moment generating function of privacy loss random variable Wang et al. (2019), Rényi DP, and subsampled-RDP-which yields tight composition for many small releases and scales well to large $n,p$ without forming $\mathbf{X}^{\top}\mathbf{X}$. (Abadi et al. (2016); Bun and Steinke (2016); Mironov (2017); Wang et al. (2019)).
@@
We favor gradient perturbation for multi-stage estimators like IVaR because it (i) composes tightly across many noisy steps using modern privacy accountants, (ii) avoids spectrum-dependent blow-ups from noising $\mathbf{X}^{\top}\mathbf{X}$ (Sheffet (2017)) and (iii) yields strong convergence rates while fitting standard training pipelines (including using minibatches, streaming, early stopping) and enabling modular, stage-wise design, which is preferable for practice (Bassily et al. (2014), Abadi et al. (2016)).
@@
Instrumental Variable Regression (IVaR) has been extensively studied in econometrics Angrist and Krueger (2001); Angrist and Pischke (2009). Classical methods such as two-stage least squares (2SLS) admit closed-form solutions but face limitations in modern applications: they do not scale well to high-dimensional or streaming data, cannot easily incorporate regularization, and are restricted to linear models. This has motivated optimization-based approaches, including convex–concave formulations of nonlinear IV Muandet et al. (2020), stochastic optimization methods for scalable and online estimation Della Vecchia and Basu (2023); Chen et al. (2024); Fonseca et al. (2024), and bi-level gradient descent algorithms with convergence guarantees Liang et al. (2025). Extensions to nonlinear IV include kernel-based methods Singh et al. (2019) and DeepIV Hartford et al. (2017). Despite these advances, prior work assumes unrestricted access to the data and does not provide end-to-end differential privacy guarantees, which are increasingly critical in sensitive domains such as healthcare, finance, and online platforms. To our knowledge, no existing method offers DP guarantees with finite-sample convergence rates for linear IV/2SLS that explicitly account for instrument strength, sample size, dimension, and iteration complexity.
@@ 2.1 Privacy notions
While $(\varepsilon,\delta)$-DP is the most widely used notion of privacy, it can be too coarse for analyzing iterative mechanisms, as composition accumulates $\varepsilon$ and $\delta$ linearly. In contrast, zero-concentrated differential privacy (zCDP) characterizes privacy loss through Rényi divergences, which ensures that the privacy loss random variable enjoys a sub-Gaussian concentration property. This yields two key benefits: (i) tighter composition, since zCDP parameters add under composition, and (ii) smooth conversion, since $\rho$-zCDP implies $(\varepsilon,\delta)$-DP with $\varepsilon=\rho+2\sqrt{\rho\log(1/\delta)}$; see Bun and Steinke (2016, Proposition 1.3). As a result, zCDP provides stronger and more analytically tractable guarantees than $(\varepsilon,\delta)$-DP, particularly in the analysis of iterative algorithms.
@@ 2.2 IVaR Model and Assumptions
We impose the following standard assumptions for IVaR model.
@@ 3 Algorithm and Theoretical Guarantees
where $0<\kappa(\tau)<1$ is the contraction rate. The specific definitions of $\delta(\tau),\bar{\gamma}(\tau),\underline{\gamma}(\tau),$ and $\kappa(\tau)$ are deferred to (9).
The proof of Theorem 3.1 is presented in Appendix C. We now offer several remarks regarding this theorem. In the presentation of Theorem 3.1, all constants $c_{0},c_{1},c_{2}$ and scaling factors hidden in $\lesssim$” are independent of major parameters $n,p,q,\rho_{1},\rho_{2},\tau$. These constants only depend on problem-specific parameters $\bm{\beta},\bm{\Theta},\sigma_{z},\sigma_{1},\sigma_{2}$.
@@
It can be easily shown that under Assumption 2, the sufficient condition for learning rates to guarantee monotonic convergence are $0<\eta_{GD}<2$ and $0<\alpha_{GD}<{2}/{\|\bm{\Theta}\|^{2}}$. We note that in our learning rate condition (3), we introduce $\delta(\tau)$ and $\psi(\tau)$ to account for the randomness in data. If we have infinite samples, the condition (3) becomes
@@
Comparing to $\eta_{GD}$ and $\alpha_{GD}$, notice that we have the same $\eta$ condition. However, the $\alpha$ condition is slightly tighter to control the randomness in first-stage estimates $\bm{\Theta}^{(t)}$.
@@
We note that the error rate (8) has an additional $\sqrt{p}$ factor compared to the error rate of $2SLS$ estimator $\|\hat{\bm{\beta}}-\bm{\beta}\|$ (see Lemma D.7 for the precise statement). This observation is further confirmed by simulations in Appendix G.1. We believe, a fundamentally different modification of the algorithm may be required to algorithmically match the rate of convergence of $2SLS$ estimator $\|\hat{\bm{\beta}}-\bm{\beta}\|$ exactly even in the no-privacy setting.
@@ 4 Experiments
We conduct experiments using both synthetic data and real data to validate our theoretical findings.

### v2

@@ Abstract
We study instrumental variable regression (IVaR) under differential privacy constraints. Classical IVaR methods (like two-stage least squares regression) rely on solving moment equations that directly use sensitive covariates and instruments, creating significant risks of privacy leakage and posing challenges in designing algorithms that are both statistically efficient and differentially private. We propose a noisy two-state gradient descent algorithm that ensures $\rho$-zero-concentrated differential privacy by injecting carefully calibrated noise into the gradient updates. Our analysis establishes finite-sample convergence rates for the proposed method, showing that the algorithm achieves consistency while preserving privacy. In particular, we derive precise bounds quantifying the trade-off among optimization, privacy, and sampling error. To the best of our knowledge, this is the first work to provide both privacy guarantees and provable convergence rates for instrumental variable regression in linear models. We further validate our theoretical findings with experiments on both synthetic and real datasets, demonstrating that our method offers practical accuracy-privacy trade-offs.
@@ 1 Introduction
Instrumental variable regression (IVaR) is a foundational tool in causal inference, designed to recover structural parameters when standard estimators fail due to endogeneity. In many observational settings, covariates are influenced by unobserved confounders, causing naive methods (such as the ordinary least squares (OLS) in the context of linear regression) to produce biased and inconsistent estimates. IVaR circumvents this by leveraging instruments, which are variables that are predictive of the endogenous regressors but independent of hidden confounders, to enable consistent estimation of causal effects (Hausman, 2001; Wooldridge, 2010; Angrist and Krueger, 2001). This perspective is increasingly important in machine learning, for example in recommendation systems where user exposure is confounded by prior preferences (Si et al., 2022), or in reinforcement learning where actions and rewards are jointly influenced by unobserved context (Xu et al., 2023). In such settings, IVaR provides a principled way to disentangle causal effects from spurious correlations, enabling more reliable decision making.
However, many applications of IVaR involve sensitive data, such as individual health records, financial transactions, or user interactions, where protecting privacy is of paramount importance. In such settings, releasing model estimates or even intermediate statistics can leak information about individuals in the dataset. Differential privacy (DP) (Dwork et al., 2006) provides a mathematically rigorous framework to ensure that an algorithm’s output does not reveal sensitive information about any single data point. Despite the importance of IVaR in causal inference, to the best of our knowledge, there are no prior works addressing the problem of performing IVaR under differential privacy. This gap motivates the central question of this paper:
@@
Our work focuses on answering this question in the context of linear IVaR models. To situate our contributions, we briefly review existing work on DP methods for OLS regression, with additional discussion in Section 1.1. Several predominant approaches have emerged in the literature: (i) perturbion methods, where the empirical covariance and cross-covariance matrices are privatized before solving the normal equations; (ii) consensus-based methods, including propose-test-release and exponential mechanism approaches, which directly privatize the estimator through carefully designed randomized output rules; and (iii) gradient perturbation methods, where iterative optimization algorithms are made private by clipping gradients and injecting calibrated Gaussian noise. While all three approaches ensure differential privacy, gradient perturbation combined with clipping has been shown to yield the sharpest statistical rates in OLS regression, particularly in high-dimensional and finite-sample regimes (Bassily et al., 2014; Brown et al., 2024a).
Given the centrality of IVaR in causal inference, it is natural to explore whether the aforementioned techniques can be adapted to this setting. Unlike OLS, however, IVaR is based on moment conditions involving both covariates and instruments, making it less straightforward to design private algorithms. In particular, sufficient-statistics perturbation and consensus-based methods have not been explored, and their adaptation is non-trivial due to the inherent ill-posedness of IVaR under weak instruments and the sensitivity of the moment equations. These challenges stem from the fact that the closed-form 2SLS estimator depends on nested matrix multiplications and inversions, whose sensitivities are difficult to characterize directly. To address this, we design a two-stage gradient-descent-based algorithm that enables injecting noise at the iteration level in a principled manner. This structure allows us to rigorously control sensitivity, calibrate the noise, and derive non-asymptotic utility guarantees.
Specifically, we make the following contributions in this work:
  • We introduce DP-2S-GD (Algorithm 1), the first differentially private algorithm for instrumental variable regression, built on noisy gradient descent with gradient clipping.
  • We establish finite-sample non-asymptotic error bound for DP-2S-GD (Theorem 3.1), explicitly characterizing the trade-off between optimization, privacy, and sampling error. The main technical challenge is to carefully control the interaction between privacy-induced noise and the contraction of the gradient dynamics across iterations, with the privacy guarantee ensured by Proposition 3.1.
@@
By developing differentially private algorithm for IVaR and establishing its theoretical bound, we enable researchers to perform robust causal analyses without compromising the confidentiality of individuals in the dataset.
@@ 1.1 Related work
Differential Privacy for Regression. One can group private regression methods into the following broad families. (1) Output/objective perturbation (private empirical risk minimization (ERM)): add noise to the final estimator (output perturbation) or inject a random linear/quadratic term into a strongly convex loss before optimizing (objective perturbation); these one-shot mechanisms give $(\varepsilon,\delta)$-DP guarantees and excess-risk bounds for convex ERM (Chaudhuri et al. (2011); Kifer et al. (2012); Bassily et al. (2014)). Recent refinements, e.g. Redberg et al. (2023), leverage subsampling and tighter accounting to improve accuracy. (2) Sufficient-statistics (matrix) perturbation: release noisy surrogates of ($\mathbf{X}^{\top}\mathbf{X},\mathbf{X}^{\top}\mathbf{y}$) (or related second-moment structures) and then solve the (regularized) normal equations; this route enables OLS-specific inference but can suffer under ill-conditioning because noise is injected at the Gram-matrix level (Dwork et al. (2014); Sheffet (2017)). Tsfadia et al. (2022) proposes a subsample-and-aggregate framework that can, in principle, be adapted to regression settings by privately estimating the relevant sufficient statistics on carefully selected data subsets. Further developments in this direction include Bernstein and Sheldon (2019) and  Ferrando and Sheldon (2024). (3) Exponential mechanism: privately selects an output by randomly choosing among candidates with probabilities that grow exponentially with their quality score, with parameters controlling how strongly it favors the higher-scoring options. This mechanism is frequently applied in constructing algorithm to privately select a regression model from a pool of non-private OLS fits on subsets of the data (Ramsay and Chenouri (2021), Cumings-Menon (2022), Amin et al. (2022)). (4) Gradient perturbation (DP-(S)GD): clip per-example (mini-batch or full) gradients and add Gaussian noise at each step, tracking privacy with bounded log moment generating function of privacy loss random variable Wang et al. (2019), Rényi DP, and subsampled-RDP-which yields tight composition for many small releases and scales well to large $n,p$ without forming $\mathbf{X}^{\top}\mathbf{X}$. (Abadi et al. (2016); Bun and Steinke (2016); Mironov (2017); Wang et al. (2019)).
@@
We favor gradient perturbation for multi-stage estimators like IVaR because it (i) composes tightly across many noisy steps using modern privacy accountants, (ii) avoids spectrum-dependent blow-ups from noising $\mathbf{X}^{\top}\mathbf{X}$ (Sheffet (2017)) and (iii) yields strong convergence rates while fitting standard training pipelines (including using minibatches, streaming, early stopping) and enabling modular, stage-wise design, which is preferable for practice (Bassily et al. (2014), Abadi et al. (2016)). Although there exists DP techniques for estimating gram matrices that avoid blow-ups, e.g., via carefully calibrated noise or regularization (Brown et al. (2023), Kamath et al. (2019)), the purely sufficient statistics-based pipelines require larger sample sizes (polynomial to condition number) than gradient-based approaches to reach a comparable accuracy in high dimensions (Brown et al., 2024b). That said, we note that the convergence of private first-order gradient methods still depend heavily on the condition number Varshney et al. (2022); Liu et al. (2023).
@@
Instrumental Variable Regression (IVaR) has been extensively studied in econometrics (Angrist and Krueger, 2001; Angrist and Pischke, 2009). Classical methods such as two-stage least squares (2SLS) admit closed-form solutions but face limitations in modern applications: they do not scale well to high-dimensional or streaming data, cannot easily incorporate regularization, and are restricted to linear models. This has motivated optimization-based approaches, including convex–concave formulations of nonlinear IV (Muandet et al., 2020), stochastic optimization methods for scalable and online estimation (Della Vecchia and Basu, 2023; Chen et al., 2024; Fonseca et al., 2024), and bi-level gradient descent algorithms with convergence guarantees (Liang et al., 2025). Extensions to nonlinear IV include kernel-based methods (Singh et al., 2019) and DeepIV (Hartford et al., 2017). Despite these advances, prior work assumes unrestricted access to the data and does not provide end-to-end differential privacy guarantees, which are increasingly critical in sensitive domains such as healthcare, finance, and online platforms. To our knowledge, no existing method offers DP guarantees with finite-sample convergence rates for linear IV/2SLS that explicitly account for instrument strength, sample size, dimension, and iteration complexity.
@@ 2.1 Privacy notions
[…]
@@ 2.2 IVaR Model and Assumptions
Optimization problem 2 admits a closed-form solution. A classical approach to solve (2) is the two-stage least squares (2SLS) estimator; see Definition 2.3.
Definition 2.3 (2SLS estimator).
Given observational data $(\mathbf{Z},\mathbf{X},\mathbf{Y})=\{(\mathbf{z}_{i},\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$, the 2SLS estimator $\hat{\bm{\beta}}_{\textsf{2SLS}}$ is obtained through two consecutive OLS regressions:
  • i.
    First stage: Regress $\mathbf{X}$ on $\mathbf{Z}$ to obtain $\hat{\bm{\Theta}}$
    $\hat{\bm{\Theta}}=(\mathbf{Z}^{\top}\mathbf{Z})^{-1}\mathbf{Z}^{\top}\mathbf{X}.$
  • ii.
    Second stage: Regress $\mathbf{Y}$ on $\hat{\mathbf{X}}:=\mathbf{Z}\hat{\bm{\Theta}}$ to obtain $\hat{\bm{\beta}}_{\textsf{2SLS}}$:
    $\hat{\bm{\beta}}_{\textsf{2SLS}}=(\hat{\bm{\Theta}}^{\top}\mathbf{Z}^{\top}\mathbf{Z}\hat{\bm{\Theta}})^{-1}\hat{\bm{\Theta}}^{\top}\mathbf{Z}^{\top}\mathbf{Y}.$
In the following sections, we will use $\hat{\bm{\beta}}$ to denote the 2SLS estimator for simplicity. We impose the following standard assumptions for IVaR model.
@@ 3 Algorithm and Theoretical Guarantees
For any fixed $\bm{\Theta}\in\mathbb{R}^{q\times p}$ and $\bm{\beta}\in\mathbb{R}^{p}$, consider the Algorithm 1 with fixed step sizes satisfying
@@
where $0<\kappa(\tau)<1$ is the contraction rate, $\delta(\tau)>0$ is a numerically small term, and $\bar{\gamma}(\tau),\underline{\gamma}(\tau)$ are the high-probability upper/lower bounds on the eigenvalues of $\frac{\hat{\bm{\Theta}}^{\top}\mathbf{Z}^{\top}\mathbf{Z}\hat{\bm{\Theta}}}{n}$. The specific definitions of $\delta(\tau),\bar{\gamma}(\tau),\underline{\gamma}(\tau),$ and $\kappa(\tau)$ are deferred to (10).
The proof of Theorem 3.1 is presented in Appendix C. We now offer several remarks regarding this theorem. In the presentation of Theorem 3.1, all constants $c_{0},c_{1},c_{2}$ and scaling factors hidden in $\lesssim$” are independent of major parameters $n,p,q,T,\rho_{1},\rho_{2},\tau$. These constants only depend on problem-specific parameters $\bm{\beta},\bm{\Theta},\sigma_{z},\sigma_{1},\sigma_{2}$.
@@
It can be easily shown that under Assumption 2, the sufficient condition for learning rates to guarantee monotonic convergence are
$0<\eta_{GD}<2,\quad 0<\alpha_{GD}<\frac{2}{\|\bm{\Theta}\|^{2}}.$
We note that in our learning rate condition (3), we introduce $\delta(\tau)$ and $\psi(\tau)$ to account for the randomness in data. If we have infinite samples, the condition (3) becomes
@@
Comparing to $\eta_{GD}$ and $\alpha_{GD}$, notice that we have the same $\eta$ condition. However, the $\alpha$ condition is slightly tighter to control the randomness introduced by the first-stage estimates $\bm{\Theta}^{(t)}$.
Remark 3.3.
From (10), the optimal contraction rate $\kappa^{\star}(\tau)$ is achieved when the learning rates are set as
$\eta_{\text{approx}}^{\star}=\frac{2}{(1+\delta(\tau))^{2}+(1-\delta(\tau))^{2}},\quad\alpha_{\text{approx}}^{\star}=\frac{2}{\bar{\gamma}(\tau)+\underline{\gamma}(\tau)}.$ (8)
In this case, we have
$\kappa_{\bm{\beta}}^{\star}(\tau)=\frac{\bar{\gamma}(\tau)}{\bar{\gamma}(\tau)+\underline{\gamma}(\tau)},\quad\kappa_{\bm{\Theta}}^{\star}(\tau)=\frac{(1+\delta(\tau))^{2}-(1-\delta(\tau))^{2}}{(1+\delta(\tau))^{2}+(1-\delta(\tau))^{2}},\kappa^{\star}(\tau)=\max\left\{\kappa_{\bm{\beta}}^{\star}(\tau),\kappa_{\bm{\Theta}}^{\star}(\tau)\right\}.$
We emphasize that although $\eta_{\text{approx}}^{\star}$ and $\alpha_{\text{approx}}^{\star}$ minimize the contraction rate, they should be viewed as approximately optimal step sizes, as the scaling constants in the bound (7) vary with different choices of step sizes. Empirically, we also observe that the estimator’s error is fairly insensitive in a neighborhood of $\eta_{\text{approx}}^{\star},\alpha_{\text{approx}}^{\star}$. See Appendix G.1 for empirical results.
@@
[…]
@@ 4 Experiments
We conduct experiments using both synthetic data and real data to validate our theoretical findings. For all experiments, we set $\tau=5$, and step sizes $\eta=\frac{1}{(1+\delta(\tau))^{2}}$, $\alpha=\frac{2}{2\bar{\gamma}(\tau)+\underline{\gamma}(\tau)}$. As a practical guideline, $\rho=0.1$ is considered as strong privacy, $\rho=1$ is considered as moderate privacy, and $\rho=10$ is considered as weak privacy [^3] The corresponding $(\epsilon,\delta)$-DP values using the conversion formula $\epsilon=\rho+2\sqrt{\rho\log(1/\delta)}$ (with $\delta=10^{-5}$): $\rho=0.1\Leftrightarrow(\epsilon,\delta)=(2.25,10^{-5})$, $\rho=1\Leftrightarrow(\epsilon,\delta)=(7.79,10^{-5})$, and $\rho=10\Leftrightarrow(\epsilon,\delta)=(31.47,10^{-5})$..